OpenStudy (anonymous):

Solve for x under the assumption that x > 0. Enter your answer in interval notation using grouping symbols. x−24/x<10

OpenStudy (anonymous):

\[x-\frac{ 24 }{ x }<10\]

OpenStudy (shootingstar28):

Let's find the critical points of the inequality. x2−24/x=10 x2−24=10x (Multiply both sides by x) x2−24−10x=10x−10x (Subtract 10x from both sides) x2−10x−24=0 (x+2)(x−12)=0 (Factor left side of equation) x+2=0 or x−12=0 (Set factors equal to 0) x=−2 or x=12

OpenStudy (anonymous):

1) Note that x is undefined at 0 x≠0 2) Replace the inequality sign with an equal sign, so that we can solve it like a normal equation x−24x=10 3) Multiply both sides by the common denominator x x2−24=10x 4) Move all terms to one side x2−24−10x=0 5) Factor x2−24−10x 1. Ask: Which two numbers add up to -10 and multiply to -24? 2. Answer: -12 and 2 3. Rewrite the expression as (x−12)(x+2): (x−12)(x+2)=0 6) Solve for x 1. Ask: When will (x−12)(x+2) equal zero? 2. Answer: When x−12=0 or x+2=0. 3. Solve each of the 2 equations above: x=12,−2 7) From the values of x above, we have these 4 intervals to test x<−2−2<x<00<x<12x>12 8) Pick a test point for each interval For the interval x<−2: Let's pick x=−3. Then, −3−24−3<10. After simplifying, we get 5<10, which is true. Keep this interval. For the interval −2<x<0: Let's pick x=−1. Then, −1−24−1<10. After simplifying, we get 23<10, which is false. Drop this interval. For the interval 0<x<12: Let's pick x=1. Then, 1−241<10. After simplifying, we get −23<10, which is true. Keep this interval. For the interval x>12: Let's pick x=13. Then, 13−2413<10. After simplifying, we get 11.1538461539<10, which is false. Drop this interval. 9) Therefore, x<−20<x<12 Hope this helps! :D

OpenStudy (shootingstar28):

Check possible critical points. x=−2(Works in original equation) x=12(Works in original equation) Critical points: x=−2 or x=12(Makes both sides equal) x=0(Makes left denominator equal to 0) Check intervals in between critical points. (Test values in the intervals to see if they work.) x<−2(Works in original inequality) −2<x<0(Doesn't work in original inequality) 0<x<12(Works in original inequality) x>12(Doesn't work in original inequality)

OpenStudy (shootingstar28):

The answer shall be: Answer: x<−2 or 0<x<12

OpenStudy (anonymous):

thanks #ShootingStar28

OpenStudy (shootingstar28):

You are welcome :)

OpenStudy (anonymous):

can you help with another one? I think I know how to do it now, just want you to check my answer. I can put it as a new quetions, if you want

OpenStudy (anonymous):

\[x-\frac{ 15 }{ x }<-2\]

OpenStudy (anonymous):

Add \[\frac{ 15 }{ 2 }\] to both sides \[x<-2+\frac{ 15 }{ 2 }\] Then simplify \[-2+\frac{ 15 }{ 2 }\] to \[\frac{ 11 }{ 2 }\] So, \[x<\frac{ 11 }{ ? }\] Hope this helps! :)

OpenStudy (anonymous):

Oops, I meant \[x=\frac{ 11 }{ 2 }\]

OpenStudy (anonymous):

Ugh, Sorry \[x<\frac{ 11 }{ 2 }\]